Skill Set and Certifications

ari.edelkind@episec.com New York, NY 917 968 6991

Ari Edelkind_{July, 2004}

Computer Skills

General experience ratings:

Some: I have used this skill enough to feel comfortable doing so, and I am eager to learn more. I also have a significant theoretical understanding of the subject. Moderate: I have an advanced theoretical understanding of this subject, and a significant technical unserstanding. I use this skill in practice, though I try to keep a book or other documentation handy while doing so. High: I have an advanced understanding of this subject, both theoretical and technical. If the subject is an operating system, I have ported applications to it and spent time examining kernel internals. If desired, I can be expected to pass certification tests on this topic without studying. Note: When sensible, all entries are given in alphabetical order.

Operating Systems

AIX 5.2 ....................................................... Moderate BSDI (BSD/OS) 2.1 – 3.1 .......................................... High DEC Digital Unix (OSF1) 3.0 ................................. Moderate DGUX 4.20 ................................................... Moderate FreeBSD 2.1 – 5.2 ................................................. High HP-UX 10.20, 11.23 ........................................... Moderate IRIX 6.3 – 6.5.4 .................................................... High Linux (multiple distributions), kernels 1.3 – 2.6 .................... High NetBSD 1.4.3 – 1.5.2 ............................................... High OpenBSD 2.1 – 2.9 ................................................ High OpenStep 4.2 ..................................................... Some QNX 4.24, 6.1 .................................................... Some Solaris (SunOS, SPARC) 2.3 – 8 .................................... High

Other (possibly) relevant operating systems

Cisco IOS and Supervisor Software ........................... Moderate MacOS 7.5.1 – 9, X (10.3) .......................................... High Windows 3.1, 95, 98, 2000, NT 4 ................................... High

Programming

Assembly ................................................... Moderate • i386, plus small amounts of SPARC and MIPS

• Used on DGUX, FreeBSD, Linux, NetBSD, OpenBSD

• Nasm and AT&T formats C (ANSI and traditional) ......................................... High

• network sockets

• kernel programming

• pty/tty manipulation

• SVR4/BSD compatibility

• low-level data structure manipulation

• dynamic libraries, library call wrapping HTML ........................................................... High JavaScript ................................................... Moderate Perl 5 ............................................................ High PHP ............................................................. High Shell scripting ................................................... High

• sh, csh, ksh SQL ......................................................... Moderate TCL ......................................................... Moderate

Security

Auditing ......................................................... High • Code auditing (C, Perl, PHP, Shell scripts)

• Network and host auditing chroot() .......................................................... High

• functionality

• implementation

• breaking Cryptography ............................................... Moderate

• Cryptographic theory ................................... High • Cryptographic libraries ............................. Moderate

(ssl, rsa, des, 3des, blowfish, md5, rijndael)

• Incorporation into programs ........................ Moderate • Cryptographic software (gnupg, pgp, kerberos, openssl)

Exploit creation (from scratch) .............................. Moderate

• Debugging, process tracing, and library call wrapping to find exploitable problems in closed-source software

• Network sniffing and monitoring

• Use of Assembly, C, and Perl to write exploits for pedagogical

demonstrations

Firewalls ........................................................ High • Firewall theory

• Firewall design

• Transparent firewalls (masquerading, NAT) Host security .................................................... High

• Securing network services

• Setuid program evaluation and revocation

• Temporary (world-writable) directory issues

• Host-based filtering (ipf, netfilter, pf)

• DSO wrapping to enhance application security measures Intrusion detection .......................................... Moderate

• IDS theory .............................................. High • IDS circumvention theory .......................... Moderate

Penetration testing .......................................... Moderate Quotas (disk, memory, cpu) ...................................... High

General service setup/administration

Including, not limited to …

Backup services (amanda, burt, dump …) ....................... High FTP servers (ncftpd, proftpd, vsftpd, wu-ftpd …) ................ High IMAP servers (courier, uw-imap) ............................ Moderate Inet services (inetd, xinetd, tcpserver) ............................ High Mail servers (sendmail, qmail) ................................... High Nameservers (bind, djbdns) ...................................... High Network file systems (nfs, samba, netatalk) ...................... High Phone database servers ..................................... Moderate POP3 servers (qpopper, popa3d, qmail-pop3d) ................... High Printing services (lprng) ......................................... High

• BSD and SVR4 line printing services ................ Moderate Shell services (ssh, openssh, telnet, evil r-commands) ............ High Single sign-on (cryptocard) ................................. Moderate SQL servers (mysql, postgresql, sqlite) ........................... High User management services ....................................... High

• Centralized management (nis, nis+)

• Distributed management (custom scripts) Web servers (apache, ncsa, publicfile) ............................ High

• netscape enterprise ..................................... Some Web server modules (php, mod perl, mod ssl) ................... High Web-based services .............................................. High

• e-mail (imp, ilohamail)

• inventory system (custom)

• password management (custom)

• project tracking (keystone)

Certifications

It is this individual’s opinion that certifications do not accurately portray one’s knowledge; therefore, certifications are made available upon request.